A jar contains 4 red balls and 6 blue balls. If three balls are drawn in sequence from the jar, what is the probability that two are red and one is blue?
Answer #1

number of ways to draw 3 balls from 10 balls =10C3 =120

number of ways to draw 2 red and 1 blue ball=4C2*6C1 =6*6=36

hence probability that two are red and one is blue =36/120=3/10=0.3